OpenStudy (kanwal32):

i have solved this question but I wanted to share this with OS it took me 1/2 hour Consider two different GP with their sum S1and S2 S1=a+ar+ar^2.......infinite=1 S2=b+bR+bR^2+bR^3.....infinite=1 S1=S2=1,ar=bR and ar^2=1/8

OpenStudy (kanwal32):

ques Sum of their first terms=? Common ratio of first G.P=? Common Ration of Second GP=?

OpenStudy (kanwal32):

note aris not equal to br

ganeshie8 (ganeshie8):

looks interesting...

OpenStudy (kanwal32):

@ganeshie8 is solved it ans1)1 Ans2)\[1-\sqrt5/4\] Ans3)\[3-\sqrt5/4\]

OpenStudy (kanwal32):

good questions

ganeshie8 (ganeshie8):

guess i would start here : \[\large \dfrac{a}{1-r} = \dfrac{b}{1-R} = 1\]

OpenStudy (kanwal32):

so u'll get equation a=1-r and substitute in ar^2

OpenStudy (kanwal32):

check that u'll get relation R+r=1

ganeshie8 (ganeshie8):

\(\large ar^2=1/8\) \(\large (1-r)r^2=1/8\) like this ?

OpenStudy (kanwal32):

in the equtaion u have wriitten

OpenStudy (kanwal32):

ya

ganeshie8 (ganeshie8):

how did u get R+r = 1 ?

OpenStudy (kanwal32):

ar=bR

OpenStudy (kanwal32):

R-R^2=r-r^2

OpenStudy (kanwal32):

i think u'll get using a^2-b^2

ganeshie8 (ganeshie8):

not sure, let me grab paper and pen lol this looks bit involved...

OpenStudy (anonymous):

*** ( Doing this as i am doing sequences and series and so that i get notified)

OpenStudy (ikram002p):

S1=a+ar+ar^2.......infinite=1 sub a+ar from both sides S1-(a+ar)=1-(a+ar) S1-(a+ar)=a+ar+ar^2.......infinite= 1-(a+ar) S1-(a+ar) = \dfrac{ar^2}{1-r} = 1-a-ar mmm does this work ?

OpenStudy (ikram002p):

\( S_1-(a+ar) = \dfrac{ar^2}{1-r} = 1-a-ar\)